# Generating Shrinkers by Mean Curvature Flow

### v2

@@ 4. Platonic and Regular Polygon Shrinkers
We can assume that each $M_{n}$ is nonfattening. For let $f_{n}:\mathbf{R}^{3}\to\mathbf{R}$ be a smooth function such that $f_{n}^{-1}(0)=M_{n}$, such that 00 is not a critical value of $f$, and such that $f>0$ at all points of $Q$. Note that there are $\epsilon_{n}\uparrow 0$ such that if $\epsilon_{n}<s_{n}<0$, then the surfaces $M_{n}^{\prime}=f^{-1}(s_{n})$ will converge smoothly to $M$. We can choose the $s_{n}$ so that each $M_{n}^{\prime}$ is nonfattening (since $f^{-1}(s)$ fattens for at most countable set of $s$.) Finally, note that $M_{n}^{\prime}(t)$ never intersects $Q$, since $M_{n}^{\prime}$ lies in the region enclosed by $M_{n}$. Thus we can assume $M_{n}$ is nonfattening (by replacing $M_{n}$ by $M_{n}^{\prime}$, if necessary.)
@@
converges smoothly and with multiplicity $1$ to a cylinder $\Sigma$ whose axis is the line $\mathbf{R}p:=\{sp:s\in\mathbf{R}\}$. Let $A$ be an Angenent torus that lies in the non-simply connected component of $\Sigma$ and that collapses to the origin.
Then, for all $t<T_{Q}$ sufficiently close to $T_{Q}$, $A$ and 00 will lie in different connected components of the surface (3). Thus the Angenent torus
$\tilde{A}:=(T_{Q}-t)A+p$
and the point $p$ will lie in different components of $M(t)$.
Fix such a time $t$ that is regular for $M(\cdot)$ and also for each of the $M_{n}(\cdot)$.
Then $M_{n}(t)$ converges smoothly to $M(t)$, so, for all sufficiently large $n$, $\tilde{A}$ and $p$ lie in different components of $M_{n}(t)$.
@@
Suppose $\mathscr{F}$ is a family of compact, connected, smoothly embedded, $G$-invariant surfaces in $\mathbf{R}^{3}\setminus Q$. Suppose also that $\mathscr{F}$ is connected with respect to smooth convergence, and that all the $M\in\mathscr{F}$ lie in some large ball $B(0,R)$. Let $\mathscr{S}$ be the subfamily of $M\in\mathscr{F}$ for which $T_{Q}(M)<\infty$. If $\mathscr{S}$ and $\mathscr{S}^{c}:=\mathscr{F}\setminus\mathscr{S}$ are both nonempty, then there is an $M\in\mathscr{S}$ for which
@@
Under mean curvature flow, the sphere $\partial B(0,R)$ shrinks to a point in time $(R/2)^{1/2}$. Thus, if $M\in\mathscr{F}$, then $M(t)$ is empty for all $t>(R/2)^{1/2}$. We claim that $\mathscr{S}$ is a closed subset of $\mathscr{F}$. To see that, suppose that $M_{n}\in\mathscr{S}$ converges to $M\in\mathscr{F}$. By hypothesis, there exist $t_{n}$ and $p_{n}$ such that
@@
Since $p_{n}\in B(0,R)$ and since $t_{n}\leq(R/2)^{1/2}$, we can assume that $p_{n}$ and $t_{n}$ converge so to finite limits $p$ and $t$. Thus
@@
    For each $s\in(0,1)$, each component of $\{f\geq s\}$ is geodesically convex.
@@
$h:(0,\infty)\to[0,\infty)$
be a smooth function such that be a smooth function such that
@@
$h(r)=(\log r)^{2}.$
Extend $h$ to the interval to $[0,\infty)$ by setting $h(0)=\infty$.
Now define
$\Phi:\{p:|p|\in[1/2,2]\}\to[0,\infty]$
by
@@
Now for $s\in[0,\infty)$, let $M^{s}$ be the set of $p$ such that $p\in[1/2,2]$ and such that
$\Phi(p)=s.$
Now $M^{0}=E$, which vanishes instantly under level set flow, so $T_{Q}(M^{0})=\infty$ and therefore $T_{Q}(M^{s})=\infty$ for all sufficiently small $s$.
Note that
$\{\Phi<\infty\}=\{p:1/2<|p|<1\}\setminus Q.$
Thus, for all sufficiently large $s$, the region $\Phi<s$ enclosed by $M^{s}$ contains an Angenent torus centered at a point $p\in C\subset Q$. Thus $T_{Q}(M^{s})<\infty$ by Lemma 12. Now fix a $c$ such that $T_{Q}(M_{c})<\infty$. Then the family
=$\mathscr{F}=\{M^{s}:0<s\leq c\}$
@@
Now consider the when $Q$ consists of the rays
$\{(r\cos\theta,r\sin\theta,0):r>0\}$
such that $\theta$ is multiple of $2\pi/(g+1)$. Thus
$Q=\{(r\cos\theta,r\sin\theta,0):r\geq 0,\,\sin\left(\frac{g+1}{2}\theta\right)=0\}.$
@@
$f(r\cos\theta,r\sin\theta,z)=u\left(\cos\left(\frac{g+1}{2}\theta\right)^{2}\right),$
where $u(s)=\frac{s}{1-s}$.
The rest of the proof is exactly as in the Platonic solid case. ∎
@@ 5. A Topological Lemma
In §4, we produced shrinkers that are disjoint from certain collections of rays from the origin. In the remainder of the paper, we will produce shrinkers that contain certain collections of rays from the origin. This section proves basic facts about the topology of such surfaces.
Let $g\geq 1$ be an integer. Let $Q=Q_{g}$ be the union of the lines
@@
such that $\theta$ is a multiple of $\pi/(g+1)$. For $\theta\in\mathbf{R}$, let $P_{\theta}$ be the vertical plane spanned by $(\cos\theta,\sin\theta,0)$ and $\mathbf{e}_{3}$. Let $G=G_{g}$ be the group of isometries of $\mathbf{R}^{3}$ consisting of:
@@
    Rotations about $Z$ by multiples of $\pi/(g+1)$.
@@
Then $M$ has genus 00, $g$, or $2g$.
Let $c$ be the number of boundary components $C$ of $\partial M$ such that $C$ lies in $\{z>0\}$ and such that $C$ winds once around $Z$. Let $d$ be the number of points in
@@
By symmetry, the number of points of $M^{\prime}\cap Z$ is $12k+$ 1. Let $M^{\prime\prime}$ be the quotient of $M^{\prime}$ under the group consisting of rotations about $Z$ by angles that are multiples of $2\pi/(g+1)$. Then
@@
Thus, multiplying (7) by $-1$ gives
@@
Since $\operatorname{genus}(M^{\prime})\leq 2g$, we see that $(M^{\prime\prime})^{+}$ has genus 00, that $k$ is 00, $1$, or $2$, and thus that
@@
For $R>0$, let $M_{R}$ be the connected component of $M\cap B(0,R)$ that contains the origin. Since $M$ is a translator, it is connected. Thus
@@
Let $M$ be as in Corollary 23. Suppose also that $M\setminus\{z=0\}$ has genus 00.
@@
Let $\Gamma(r)=\operatorname{genus}(M\cap B(0,r))$. Trivially, $\Gamma$ is an increasing function of $r$, $\Gamma(r)$ is 00 for all sufficiently small $r$ and Let $M^{\prime}_{r}$ be the connected component of $M\cap B(0,r)$ containing the origin. The portion of $M\cap B(0,R)$ not in $M^{\prime}_{r}$ lies in
@@
Now suppose that $M$ has genus $g$. By Lemma 22, The right hand side of (9) is either 00 or $g$. Thus there is an $r$ such that
@@
$=0$
This completes the proof when $M$ has genus $g$. The case when $M$ has genus $2g$ is essentially the same; the function $\Gamma(\cdot)$ only takes the values $g$ and $2g$. We can let $r$ be the largest $r$ for which $\Gamma(r)<g$ and $R$ be the largest $R$ for which $\Gamma(r)<2g$. ∎
@@
Except where otherwise specified, $N$ will refer the manifold $N$ with the Riemannian metric $\gamma:=\gamma_{1}$.
@@
If $g$ is a positive integer, we let $Q=Q_{g}$ be set of horizontal lines and $G=G_{g}$ be the group of isometries of $\mathbf{R}^{3}$ that were defined at the beginning of Section 5. Note that each $\sigma\in G$ extends to an isometry of $N:=\mathbf{R}^{3}\cup Z_{\infty}$.
@@
Most of this section is dedicated to establishing that if $M\in\mathscr{M}_{g}$, then the evolving surface $M(t)$ remains in $\mathscr{M}_{g}$ for all regular times $t$ in the interval $[0,T_{\textnormal{pos}}(M)]$. See Theorem 31. We also analyze the behavior of $M(t)$ as $t$ approaches $T_{\textnormal{pos}}(M)$. Specifically, we demonstrate that if $T_{\textnormal{pos}}(M)$ is finite, then the origin is a singular point at time $T_{\textnormal{pos}}(M)$. Additionally, we show that each shrinker $\Sigma$ at $(O,T_{\textnormal{pos}})$ has genus $g$ or $2g$, and that its intersection with $\{z=0\}$ is $Q$. See Lemma 34 and Theorem 35.
@@
Furthermore,
$\operatorname{genus}(M^{\prime})=2\operatorname{genus}(M^{\prime}\cap U),$
where $U=\{(x,y,z):(x^{2}+y^{2})^{1/2}<2\}.$ Now $M^{\prime}\cap\overline{U}$ has exactly one boundary component, namely the circle $C$. Thus, by Lemma 22,
@@
so $M^{\prime}$ has genus $2dg$. ∎
@@
where $\pi$ is the projection $\pi(x,y,z)=(x,y,0)$ and where $\operatorname{dist}$ is geodesic distance with repect to the metric $\gamma$. We let $\operatorname{Cone}^{\prime}(\Lambda)$ be the set of all $p$ satisfying (13) Thus $\operatorname{Cone}(\Lambda)$ is the portion of $\operatorname{Cone}^{\prime}(\Lambda)$ in the set (12). Conversely, $\operatorname{Cone}^{\prime}(\Lambda)$ consists of $\operatorname{Cone}(\Lambda)$ together with its images under $\rho_{X}$, $\rho_{C}$, and $\rho_{C}\circ\rho_{X}$.
Note that if $\Lambda$ is sufficiently large, then $M\subset\operatorname{Cone}^{\prime}(\Lambda)$.
@@
Also, if $M\in\operatorname{Cone}^{\prime}(\Lambda)$, then $M(t)\in\operatorname{Cone}^{\prime}(\Lambda)$ for all $t\in[0,\operatorname{T_{\textnormal{fat}}}]$.
@@
(14) $\Gamma$ $:=\partial\,\Sigma_{0}$
@@
Figure 4. The curve $\Gamma$ in the case $g=3$.
We prove the lemma by elliptic regularization. We work in $N\times\mathbf{R}$. We let $\widetilde{G}$ (and therefore $\widetilde{G}^{+}$) act on $N\times\mathbf{R}$ by acting on the first factor.
Let $w:N\times\mathbf{R}\to\mathbf{R}$ be the projection onto the second factor. For $\lambda\geq 0$, let $g_{\lambda}$ be $e^{-(2/3)\lambda w}$ times the product metric on $N\times\mathbf{R}$.
@@
Standard cut-and-paste arguments (as in [morgan]*Lemma 7.3) show that $A_{\lambda}$ is $G^{+}$-invariant. Alternatively, if one does not wish to invoke those arguments, one can simply let $A_{\lambda}$ minimize $g_{\lambda}$-area in the class of $G^{+}$-invariant surfaces.
@@
(15) $t\in(0,\infty)\mapsto A_{\lambda}(t):=A_{\lambda}-t\lambda\mathbf{e}_{4},$
$t\in(0,\infty)\mapsto A^{\prime}_{\lambda}(t):=A^{\prime}_{\lambda}-t\lambda\mathbf{e}_{4}$
@@
The points in $C$ are regular points of the flow at all times $t\in[0,\operatorname{T_{\textnormal{fat}}}]$.
@@
Let $\Sigma$ be a shrinker to the flow at a spacetime point $(p,t)$, where $p\in C$. By Lemma 28, $\Sigma$ (after a rotation) is contained in the set
$\{(x,y,z):|z|\leq|x|\}.$
@@
This is a special case of the following theorem (see [white-genus]): if $M$ is a smoothly embedded, compact surface in a complete $3$-manifold of bounded Ricci curvature and if $K$ is a closed set such that $K\subset M(t)$ for all $t\in[0,T_{\textnormal{pos}}]$, then for every regular time $t<T_{\textnormal{pos}}$,
@@
Recall that $\{z=0\}$ is isometric to a sphere of radius $1$, so it has area $4\pi$.
@@
If $\operatorname{T_{\textnormal{fat}}}=\infty$, then $M(t)$ converges smoothly to $\{z=0\}$ with multiplicity $1$. In particular, there is a time $T<\infty$ such that $M(t)$ is a smooth graph (i.e., such that $M(t)$ projects diffeomorphically onto $\{z=0\}$) for all $t\geq T$.
@@
Let $t_{n}\to\infty$. After passing to a subsequence, the flows
@@
converge to a static integral Brakke flow. That is, there is a stationary integral varifold $V$ such that for every $t$, $M_{n}(t)$ converges (as Radon measures) to the Radon measure associated to $V$.
Now $V$ is contained in $\operatorname{Cone}^{\prime}(\Lambda)$, so, by the maximum principle, $V$ is contained in $\{z=0\}$. Thus $V$ is $\{z=0\}$ with some integer multiplicity $k$. Since $M_{n}(t)$ contains $\widetilde{Q}$ for all $t$, we see that $\widetilde{Q}$ is in the support of $V$, and thus that
@@
Let $p$ be a point in $\widetilde{Q}\setminus\{O,O_{\infty}\}$. Then for all $T\in(0,\operatorname{T_{\textnormal{fat}}}(M)]$, $(p,T)$ is a regular point of the flow $M(\cdot)$.
@@
The origin is a regular point of the flow $M(\cdot)$ at all times $<T_{\textnormal{pos}}$. If $T_{\textnormal{pos}}<\infty$, then the origin is singular at time $T_{\textnormal{pos}}$, and if $\Sigma$ is a shrinker at $(O,T_{\textnormal{pos}})$, then $\Sigma$ has genus $>0$.
@@
Suppose that $T_{\textnormal{pos}}<\infty$ and that $\Sigma$ is a shrinker at the origin time $T_{\textnormal{pos}}$. Then $\Sigma\cap\{z=0\}=Q$. Furthermore,
@@
converges smoothly to $\Sigma$. Since $\widetilde{Q}\subset M(t)$ for each $t$, $Q\subset\Sigma$. Since $(0,T)$ is not a regular point (by Lemma 34), $\Sigma$ is not a plane. Thus, as $\Sigma$ and $\{z=0\}$ are both minimal with respect to the shrinker metric, $\Sigma\cap\{z=0\}$ is a network of curves, the vertices of which are the points of tangency. Hence if $\Sigma\cap\{z=0\}$ contained a point not in $Q$, it would also contain such point $q$ where $\Sigma$ and $\{z=0\}$ intersect transversely. But then $q$ would be a limit of points where $S_{i}$ and $\{z=0\}$ intersect, which is impossible since $M(t_{i})\cap\{z=0\}$ is $\widetilde{Q}$.

### v3

@@ 4. Platonic and Regular Polygon Shrinkers
We can assume that each $M_{n}$ is nonfattening. For let $f_{n}:\mathbf{R}^{3}\to\mathbf{R}$ be a smooth function such that $f_{n}^{-1}(0)=M_{n}$, such that 0 is not a critical value of $f$, and such that $f>0$ at all points of $Q$. Note that there are $\epsilon_{n}\uparrow 0$ such that if $\epsilon_{n}<s_{n}<0$, then the surfaces $M_{n}^{\prime}=f^{-1}(s_{n})$ will converge smoothly to $M$. We can choose the $s_{n}$ so that each $M_{n}^{\prime}$ is nonfattening (since $f^{-1}(s)$ fattens for at most countable set of $s$.) Finally, note that $M_{n}^{\prime}(t)$ never intersects $Q$, since $M_{n}^{\prime}$ lies in the region enclosed by $M_{n}$. Thus we can assume $M_{n}$ is nonfattening (by replacing $M_{n}$ by $M_{n}^{\prime}$, if necessary.)
@@
converges smoothly and with multiplicity $1$ to a cylinder $\Sigma$ whose axis is the line $\mathbf{R}p:=\{sp:s\in\mathbf{R}\}$. Let $A$ be a scaled Angenent torus that lies in the non-simply connected component of the complement of $\Sigma$, and that collapses to the origin.
Then, for all $t<T_{Q}$ sufficiently close to $T_{Q}$, $A$ and 0 will lie in different connected components of the complement of the surface (6). Fix such a time $t$ that is regular for flow $M(\cdot)$ and also for each of the flows $M_{n}(\cdot)$. Note that the scaled Angenent torus
(7) $\tilde{A}:=(T_{Q}-t)^{1/2}A+p$
and the point $p$ lie in different components of the complement of $M(t)$ in $\mathbf{R}^{3}$.
(The reader may notice that the right-hand side of (7) would not be a mean curvature flow if $t$ were varying. But that is irrelevant: we have fixed a time $t$ and thus are defining a single surface $\tilde{A}$, not a one-parameter family of surfaces.)
By the smooth convergence of $M_{n}(t)$ to $M(t)$, $\tilde{A}$ and $p$ lie in different components of $M_{n}(t){{}^{c}}$ for all sufficiently large $n$.
@@
Suppose $\mathscr{F}$ is a family of compact, connected, smoothly embedded, $G$-invariant surfaces in $\mathbf{R}^{3}\setminus Q$ of genus $g$. Suppose also that $\mathscr{F}$ is connected with respect to smooth convergence, and that all the $M\in\mathscr{F}$ lie in some large ball $B(0,R)$. Let $\mathscr{S}$ be the subfamily of $M\in\mathscr{F}$ for which $T_{Q}(M)<\infty$. In other words, let $\mathscr{S}=\mathscr{F}\cap\mathscr{A}$. (See Definition 14.) If $\mathscr{S}$ and $\mathscr{S}^{c}:=\mathscr{F}\setminus\mathscr{S}$ are both nonempty, then there is an $M\in\mathscr{S}$ for which
@@
Under mean curvature flow, the sphere $\partial B(0,R)$ shrinks to a point in time $R^{2}/4$. Thus, if $M\in\mathscr{F}$, then $M(t)$ is empty for all $t>R^{2}/4$. We claim that $\mathscr{S}$ is a closed subset of $\mathscr{F}$. To see that, suppose that $M_{n}\in\mathscr{S}$ converges to $M\in\mathscr{F}$. By hypothesis, there exist $t_{n}$ and $p_{n}$ such that
@@
Since $p_{n}\in B(0,R)$ and since $t_{n}\leq R^{2}/4$, we can assume that $p_{n}$ and $t_{n}$ converge so to finite limits $p$ and $t$. Thus
@@
The reader may regard Theorem 24 as obviously true. For completeness, we give a proof.
@@
    For each $s\in(0,\infty)$, each component of $\{f\geq s\}$ is geodesically convex. (There is one component in each face of $U$.)
@@
$h:(1/2,2)\to[0,\infty)$
be a smooth function such that
@@
$h(r)=\left(\log\frac{(2r-1)}{2-r}\right)^{2}.$
Extend $h$ to the closed interval by setting $h(1/2)=h(2)=\infty$.
Now let
${\bf W}:=\{p\in\mathbf{R}^{3}:1/2\leq|p|\leq 2\},$
and define $\Phi:{\bf W}\to[0,\infty]$ by
@@
For $s\in[0,\infty]$, let $M^{s}:=\Phi^{-1}(s)$. By construction, $\Phi$ is $G$-invariant and has nonvanishing gradient on
$\operatorname{interior}({\bf W})\setminus(Q\cup E)=\bigcup_{0<s<\infty}M^{s}.$
Hence the surfaces $M^{s}$ are $G$-invariant, smooth, and vary smoothly for $s\in(0,\infty)$. Also, for $s>0$ small, the genus of $M^{s}$ is the same as the genus of the boundary of a tubular neighborhood of $E=M^{0}$, which is equal to $g$. It follows that the genus of $M^{s}$ equals $g$ for all $s\in(0,\infty)$.
Now, $M^{0}=E$ vanishes instantly under level set flow, so $T_{Q}(M^{0})=\infty$ and therefore $T_{Q}(M^{s})=\infty$ for all sufficiently small $s$.
Moreover, as $s\rightarrow\infty$,
[…]
Thus, for all sufficiently large $s$, the region $\{\Phi<s\}\subset({\bf W}\setminus Q)$ enclosed by $M^{s}$ contains a suitably scaled Angenent torus centered at a point $p\in{\bf W}\cap Q$. It follows from Lemma 13 that $T_{Q}(M^{s})<\infty$. Therefore, the family
$\mathscr{F}=\{M^{s}:0<s<\infty\}$
@@
Now consider the case when $Q$ consists of the rays
$\{(r\cos\theta,r\sin\theta,0):r\geq 0\}$
@@ 5. A Topological Lemma
In §4, we produced shrinkers that are disjoint from certain collections of rays from the origin. In the remainder of the paper, we will produce shrinkers that contain certain collections of lines through the origin. This section proves basic facts about the topology of such surfaces.
Let $g\geq 1$ be an integer. Let $Q=Q_{g}$ be the union of the $g+1$ lines
@@
such that $\theta$ is a multiple of $\pi/(g+1)$. For $\theta\in\mathbf{R}$, let $P_{\theta}$ be the vertical plane spanned by $(\cos\theta,\sin\theta,0)$ and $\mathbf{e}_{3}$. Let $G=G_{g}$ be the group of isometries of $\mathbf{R}^{3}$ generated by:
@@
    Rotations about $Z$ by multiples of $2\pi/(g+1)$.
@@
(Equivalently, $G_{g}$ is the group of symmetries of the surface given in cylindrical coordinates by
$z=r^{g+1}\sin((g+1)\theta).$
The group $G_{g}$ is also the group of symmetries of the Costa-Hoffman-Meeks surface of genus $g$.)
@@
Then $M$ has genus 0, $g$, or $2g$.
Let $c$ be the number of components of $\partial M$ that lie in $\{z>0\}$ and that wind once around $Z$. Let $d$ be the number of points in
@@
By symmetry, the number of points of $M^{\prime}\cap Z$ is $2k+1$. Let $M^{\prime\prime}$ be the quotient of $M^{\prime}$ under the group consisting of rotations about $Z$ by angles that are multiples of $2\pi/(g+1)$. Then
@@
Thus, multiplying (11) by 1-1 gives
@@
Since $\operatorname{genus}(M^{\prime})\leq 2g$, we see that $(M^{\prime\prime})^{+}$ has genus 0, that $k$ is 0, 11, or $2$, and thus that
@@
For $R>0$, let $M_{R}$ be the connected component of $M\cap B(0,R)$ that contains the origin. Since $M$ is a shrinker, it is connected. Thus
@@
Let $M$ be as in Corollary 27. Suppose also that $M\setminus\{z=0\}$ has genus 0.
@@
Let $\Gamma(r)=\operatorname{genus}(M\cap B(0,r))$. Trivially, $\Gamma$ is an increasing function of $r$, and $\Gamma(r)=0$ for all sufficiently small $r$. Let $M^{\prime}_{r}$ be the connected component of $M\cap B(0,r)$ containing the origin. The portion of $M\cap B(0,R)$ not in $M^{\prime}_{r}$ lies in
@@
Now suppose that $M$ has genus $g$. By Lemma 25, The right hand side of (13) is either 0 or $g$. Thus there is an $r$ such that
@@
$=0,$
[…]
@@
Except where otherwise specified, $N$ will refer the manifold $N$ with the Riemannian metric $\gamma:=\gamma_{1}$. One easily checks that
(14) $\gamma_{ij}\leq\delta_{ij}$
at all points.
@@
If $g$ is a positive integer, we let $Q=Q_{g}$ be the set of $g+1$ horizontal lines and we let $G=G_{g}$ be the group of isometries of $\mathbf{R}^{3}$ that were defined at the beginning of Section 5. Note that each $\sigma\in G$ extends to an isometry of $N:=\mathbf{R}^{3}\cup Z_{\infty}$.
@@
Most of this section is dedicated to establishing that if $M\in\mathscr{M}_{g}$, then the evolving surface $M(t)$ remains in $\mathscr{M}_{g}$ for all regular times $t$ in the interval $0,T_{\textnormal{pos}}(M)]$. See Theorem 35. We also analyze the behavior of $M(t)$ as $t$ approaches $T_{\textnormal{pos}}(M)$. Specifically, we demonstrate that if $T_{\textnormal{pos}}(M)$ is finite, then the origin is a singular point at time $T_{\textnormal{pos}}(M)$. Additionally, we show that each shrinker $\Sigma$ at $(O,T_{\textnormal{pos}})$ has genus $g$ or $2g$, and that its intersection with $\{z=0\}$ is $Q$. See Lemma 38 and Theorem 39.
@@
Let $U=\{(x,y,z):(x^{2}+y^{2})^{1/2}<2\}$, and note that $\overline{U}=U\cup\operatorname{Cyl}$, where $\operatorname{Cyl}$ is the cylinder defined in (16). Now $M^{\prime}\cap\overline{U}$ has exactly one boundary component, namely the circle $C$ $\subset\{z=0\}$. Thus, by Lemma 25,
@@
But $M^{\prime}\setminus C$ is the disjoint union of $M^{\prime}\cap{U}$ and $\rho_{C}(M^{\prime}\cap{U})$. Hence
$\operatorname{genus}(M^{\prime})=2\operatorname{genus}(M^{\prime}\cap U).$
From the three displayed equations above, we conclude that
$\operatorname{genus}(M)=\operatorname{genus}(M^{\prime})=2dg.$
@@
where $\pi$ is the projection $\pi(x,y,z)=(x,y,0)$ and where $\operatorname{dist}$ is geodesic distance with repect to the metric $\gamma$. We let $\operatorname{Cone}^{\prime}(\Lambda)$ be the set of all $p$ satisfying (18). Thus $\operatorname{Cone}(\Lambda)$ is the portion of $\operatorname{Cone}^{\prime}(\Lambda)$ in the set (17). Conversely, $\operatorname{Cone}^{\prime}(\Lambda)$ consists of $\operatorname{Cone}(\Lambda)$ together with its images under $\rho_{X}$, $\rho_{C}$, and $\rho_{C}\circ\rho_{X}$.
Note that if $\Lambda$ is sufficiently large, then
(19) $M\subset\operatorname{Cone}^{\prime}(\Lambda).$
@@
Also, if $M\in\operatorname{Cone}^{\prime}(\Lambda)$, then $M(t)\in\operatorname{Cone}^{\prime}(\Lambda)$ for all $t\in[0,T_{\textnormal{pos}}]$.
@@
(20) $\Gamma$ $=\partial\,\Sigma_{0}$
@@
Figure 4. The curve $\Gamma$ in the case $g=4$.
We prove the lemma by elliptic regularization. We work in $N\times\mathbf{R}$. We let $\widetilde{G}$ act on $N\times\mathbf{R}$ by acting on the first factor. Recall that $\widetilde{G}^{+}$ is the subgroup of $\widetilde{G}$ that fixes the positive $z$-axis of $N$. Let $w:N\times\mathbf{R}\to\mathbf{R}$ be the projection onto the second factor. For $\lambda\geq 0$, let $g_{\lambda}$ be $e^{-(2/3)\lambda w}$ times the product metric on $N\times\mathbf{R}$.
@@
Standard cut-and-paste arguments (as in [morgan]*Lemma 7.3) show that $A_{\lambda}$ is $\widetilde{G}^{+}-$invariant. Alternatively, if one does not wish to invoke those arguments, one can simply let $A_{\lambda}$ minimize $g_{\lambda}$-area in the class of $\widetilde{G}^{+}-$invariant surfaces.
@@
(21) $t\in(0,\infty)\mapsto A_{\lambda}(t)=A_{\lambda}-t\lambda\mathbf{e}_{4},$
$t\in(0,\infty)\mapsto A^{\prime}_{\lambda}(t)=A^{\prime}_{\lambda}-t\lambda\mathbf{e}_{4}$
@@
The points in $C$ are regular points of the flow at all times $t\in[0,T_{\textnormal{pos}}]$.
@@
Let $\Sigma$ be a shrinker to the flow at a spacetime point $(p,t)$, where $p\in C$. By (19) and Lemma 32, $\Sigma$ (after a rotation) is contained in the set
$\{(x,y,z):|z|\leq\Lambda|x|\}.$
@@
This is a special case of the following fact (see [white-genus]*Corollary 22): if $M$ is a smoothly embedded, compact surface, in a complete $3$-manifold NN of bounded Ricci curvature, if $K\subset N$ is the union of finitely many smooth, simple closed curves, no two of which meet tangentially, and if $K\subset\operatorname{Reg}M(t)$ for all $t\in[0,T_{\textnormal{pos}})$, then for every regular time $t<T_{\textnormal{pos}}$,
@@
Suppose $M\in\mathscr{M}_{g}$. If $T_{\textnormal{pos}}=\infty$, then $M(t)$ converges smoothly to $\{z=0\}$ with multiplicity $1$. In particular, there is a time $T<\infty$ such that $M(t)$ is a smooth graph (i.e., such that $M(t)$ projects diffeomorphically onto $\{z=0\}$) for all $t\geq T$.
@@
Since the area of $M(t)$ is a positive, decreasing function of $t$, it has a limit $A$ as $t\to\infty$. Let $t_{n}\to\infty$. After passing to a subsequence, the flows
@@
converge to an integral Brakke flow $t\in\mathbf{R}\mapsto M^{\prime}(t)$. Note that the area of $M^{\prime}(t)$ is $A$ for every $t$. Since the area is independent of $t$, the flow is static: there is a stationary integral varifold $V$ such that $M^{\prime}(t)$ is the Radon measure associated to $V$ for $t$.
Now $V$ is contained in $\operatorname{Cone}^{\prime}(\Lambda)$ by (19) and Lemma 32, so the support of $V$ is compact. In particular, $z(\cdot)$ attains a maximum $z_{\textnormal{max}}$ on the support of $V$. It follows from the maximum principle ([sw], for example) that $\operatorname{spt}(V)$ contains all of $\{z=z_{\textnormal{max}}\}$. Since $V$ is contained in $\operatorname{Cone}^{\prime}(\Lambda)$, it follows that $z_{\textnormal{max}}=0$. Likewise, the minimum of $z(\cdot)$ on $\operatorname{spt}(V)$ is 0. Thus $V$ is $\{z=0\}$ with some integer multiplicity $k$. Since $M_{n}(t)$ contains $\widetilde{Q}$ for all $t$, we see that $\widetilde{Q}$ is in the support of $V$, and thus that
@@
Suppose $M\in\mathscr{M}_{g}$. Let $p$ be a point in $\widetilde{Q}\setminus\{O,O_{\infty}\}$. Then for all $T\in(0,T_{\textnormal{pos}}(M)]$, $(p,T)$ is a regular point of the flow $M(\cdot)$.
@@
The origin is a regular point of the flow $M(\cdot)$ at all times $T_{\textnormal{pos}}$. If $T_{\textnormal{pos}}<\infty$, then the origin is singular at time $T_{\textnormal{pos}}$, and if $\Sigma$ is a shrinker at $(O,T_{\textnormal{pos}})$, then $\Sigma$ has genus >0>0.
@@
Suppose that $M\in\mathscr{M}_{g}$ and that $T_{\textnormal{pos}}<\infty$. Let $\Sigma$ be a shrinker at the origin at time $T_{\textnormal{pos}}$. Then $\Sigma\cap\{z=0\}=Q$. Furthermore,
@@
converges smoothly to $\Sigma$. Since $\widetilde{Q}\subset M(t)$ for each $t$, $Q\subset\Sigma$. Since $(O,T)$ is not a regular point (by Lemma 38), $\Sigma$ is not a plane. Thus, as $\Sigma$ and $\{z=0\}$ are both minimal with respect to the shrinker metric, $\Sigma\cap\{z=0\}$ is a network of curves, the vertices of which are the points of tangency. Hence if $\Sigma\cap\{z=0\}$ contained a point not in $Q$, it would also contain such point $q$ where $\Sigma$ and $\{z=0\}$ intersect transversely. But then $q$ would be a limit of points where $S_{i}$ and $\{z=0\}$ intersect, which is impossible since $M(t_{i})\cap\{z=0\}$ is $\widetilde{Q}$.